OpenStudy (anonymous):

The length of the hypotenuse of a 30°-60°-90° triangle is 12. Find the perimeter.

OpenStudy (mertsj):



OpenStudy (kinggeorge):

Since it's a 30-60-90 triangle, you know the sides are proportional to 1, 2, and \(\sqrt3\). Where 2 is the length of the hypotenuse. So if you apply the same ratios, you find the perimeter should be \(6+12+\sqrt{18}\)

OpenStudy (kinggeorge):

Excuse me, it shouldn't be \(\sqrt{18}\) it should be \(\sqrt{108} = 6\sqrt{3}\)

OpenStudy (mertsj):

In a 30-60-90- triangle, the short leg is 1/2 the hypotenuse and the long leg is the short leg multiplied by sqrt3

OpenStudy (anonymous):

Can I good answer more than 1 person?

OpenStudy (kinggeorge):

You can good answer as many people as you like.

OpenStudy (anonymous):

Ok Thank you guys.